OpenStudy (anonymous):

f(2x+4)=6x+13 Find f(5-3x)+(5x-1)

OpenStudy (anonymous):

Let 2x+4=y Then x=(y-4)/2 Hence f(2x+4)=6x+13 gives f(y)=6(y-4)/2+13=3(y-4)+13 =3y-12+13=3y+1

OpenStudy (anonymous):

So f(x)=3x+1 Hence f(5-3x)=3(5-3x)+1=15-9x+1=16-9x Therefore f(5-3x)+(5x-1)=16-9x+5x-1=15-4x.@Gwin

OpenStudy (anonymous):

Can you please explain in words?

OpenStudy (anonymous):

See we are not given f(x) right??.@Gwin

OpenStudy (anonymous):

Oui

OpenStudy (anonymous):

???

OpenStudy (anonymous):

first we need to find f(x)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

read the attachment and tell me what you figured out

OpenStudy (anonymous):

You let y equal what is in the parenthesis and then plug it in to what you want to find, correct?

OpenStudy (anonymous):

yeah, we want f(y) = something Then by interchanging y by x we have f(x)

OpenStudy (anonymous):

next??

OpenStudy (anonymous):

@Gwin

OpenStudy (anonymous):

Then solve for x and plug it into f(5-3x)+f(5x-1)

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

Hi.@Gwin

OpenStudy (anonymous):

OH it was f(5-3x)+f(5x-1)?? see you typed wrong the question you posted... Yeah thats it.

OpenStudy (anonymous):

I seriously typed it wrong? @Princer_Jones

OpenStudy (anonymous):

In your question it was Find f(5-3x)+(5x-1)

OpenStudy (anonymous):

@Gwin

OpenStudy (anonymous):

Oh super.

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

So you understood how to solve this kind of problems?.@Gwin

OpenStudy (anonymous):

Sure @Princer_Jones (I use this during Math Team Class, so that's like one out of eighty problems that are hard, confusing, or both.)